When a man observed a sobriety checkpoint conducted by a police department, he saw 677 drivers were screened and 4 were arrested for driving while intoxicated. E on those results, we can estimate that P(W) =0.00591 , where w denotes the event of screening a driver and getting someone who is intoxicated. What does P (W) denote, and what is its value? What does P(W) represent? *A. P(W) denotes the probability of screening a driver and finding that he or she is not intoxicated. B. P (W) denotes the probability of driver being intoxicated. XC. P(W) denotes the probability of screening a driver and finding that he or she is intoxicated. D. P(W) denotes the probability of a driver passing through the sobriety checkpoint. P (W) - Round to five decimal places as needed.)
Answer #1

\bar W is the complement event of W.

Therefore P( \bar W) = 1 - P( W) = 1 - 0.00591 = 0.99409

So final answer is 0.99409